Setting Standards, Materials and Labour Variances

Tom Belford and Tony Sorrentino own a small business devoted to kitchen and bath granite installations. Recently, building contractors have insisted on up-front bid prices for a house rather than the cost-plus system that Tom and Tony were used to. They worry because natural flaws in the granite make it impossible to tell in advance exactly how much granite will be used on a particular job. In addition, granite can be easily broken, meaning that Tom or Tony could ruin a slab and would need to start over with a new one. Sometimes the improperly cut pieces could be used for smaller installations, sometimes not. All their accounting is done by a local chartered accounting firm headed by Charlene Davenport. Charlene listened to their concerns and suggested that it might be time to implement tighter controls by setting up a standard costing system.

Charlene reviewed the invoices pertaining to a number of Tom and Tony's previous jobs to determine the average amount of granite and glue needed per square metre. She then updated prices on both materials to reflect current conditions. The standards she developed for one square metre of counter installed were as follows:

Granite, per square metre $50.00
Glue (10 oz. @ $0.15) 1.50
Direct labour hours:
Cutting labour (0.10 hr. @ $15) 1.50
Installation labour (0.25 hr. @ $25) 6.25

These standards assumed that one seamless counter requires one sink cut (the space into which the sink will fit) as well as cutting the counter to fit the space available.

Charlene tracked the actual costs incurred by Tom and Tony for granite installation for the next six months. She found that they completed 50 jobs with an average of 32 square metres of granite installed in each one. The following information on actual amounts used and cost was gathered:

Granite purchased and used (1,640 sq. ft.) $80,848
Glue purchased and used (16,000 oz.) $2,560
Actual hours cutting labour 220
Actual hours installation labour 390

The actual wage rate for cutting and installation labour remained unchanged from the standard rate.
Answer #1

Materials price and usage variance

Granite:

Materials price variance = Actual cost * (AQ × SP)

= $80848 – (1,640 × $50) = $1152 F

Materials usage variance= (AQ – SQ)*SP

= (1,640 – 1,600)*$50 = $2,000 U

Glue:

Materials price variance = Actual cost – (AQ × SP)

= $2,560 – (16,000 × $0.15) = $160 U

Materials usage variance= (AQ – SQ) * SP

= (16,000 – 16,000)*$0.15 = 0

Labor rate and efficiency variance

Cutting Labor:

Labor rate variance = (AR – SR)*AH

= ($15 – $15)*220 = 0

Labor efficiency variance= (AH – SH) * SR

= (220 – 160)*$15 = $750 U

Installation Labor:

Labor rate variance= (AR – SR)*AH

= ($25 – $25)*390 = 0

Labor efficiency variance = (AH – SH)*SR

= (390 – 400)*$25 = $250 F